OpenStudy (anonymous):

What is the solution of the equation? 5x^2=50

OpenStudy (anonymous):

5^2 = 25 So I don't know what this is asking...

OpenStudy (anonymous):

it is x^2=50/5 which is x^2=10 and so, x=\[\sqrt{10}\] so x=+or-\[\sqrt{10}\]